# A short-baseline neutrino beam at CERN for high-precision cross-section measurements

### v1

@@ 4.2 Water-based detector
Water is the target material for Hyper-Kamiokande, making a detector based on this target an invaluable tool for achieving the SBN@CERN physics goals through precise neutrino interaction measurements in water. A widely used option, water Cherenkov detectors, offer less precise reconstruction of neutrino interactions at the GeV scale compared to liquid argon detectors, but this limitation is compensated by reduced complexity and cost.
Tests of Water Cherenkov detector performance using charged beams—mirroring the ProtoDUNE physics program—are currently ongoing at CERN with the WCTE detector [59]. WCTE is a 40-ton demonstrator of the Intermediate Water Cherenkov Detector (IWCD) [60], which is currently under construction in Japan. Water Cherenkov technology offers an ideal setting for SBN@CERN operations, even in tagged mode, as it can achieve sub-nanosecond time resolution at large scales with limited light coverage, due to the inherently faster propagation of Cherenkov light compared to liquid argon scintillation light. The WCTE/IWCD technology employs multi-PMT modules instead of the classical 20-inch PMTs used in Hyper-Kamiokande, enhancing energy linearity over a broader dynamic range and thereby improving the precision of cross-section measurements. However, the 40-ton size of WCTE is insufficient to complete the full SBN@CERN physics program in water; a dedicated detector with a fiducial volume of a few hundred tons would be required. Further enhancement of WCTE’s particle identification capability could be achieved by replacing the water target with modern water-based scintillators. This novel technology, pursued by several experimental groups, aims to overcome the limitations of Water Cherenkov detectors in medium- and large-scale experiments. It is also under consideration for the aforementioned DUNE Module of Opportunity, an initiative led by the THEIA collaborations in the US and Europe.
@@ 4.3 Reference detector parameters
The studies presented here are based on a reference liquid argon (LAr) detector with a 500-ton fiducial mass, positioned 25 m from the end of the decay tunnel. The front face measures m2, with a total length of 22.3 m. The LAr TPC is assumed to have a time resolution improved to 300 ps compared to the current ProtoDUNE, thanks to an enhanced photon detection system as planned for DUNE Phase II upgrades. The final detector optimization is still a work in progress and will depend on the chosen installation site for SBN@CERN.
@@ 6.1 The Narrow-Band Off-Axis technique
Due to the proximity of the detector to the neutrino beam, and to the large size of the detector itself, the detector surface () is exposed to a range of incoming neutrino directions. The energy of produced in and decays is correlated to the emission angle, due to the two-body decay kinematics. As a result, neutrinos interacting at different off-axis angles span different energy ranges. The off-axis technique has already been exploited by a cross-section measurement from the T2K experiment [66], and is central for the PRISM technique [67], which is planned for use in the DUNE [68] and Hyper-Kamiokande near detectors [4] as well as in the SBND experiment [20]. A key advantage of a narrow-band beam like SBN@CERN is that the off-axis neutrino energy estimation can leverage the parent meson momentum knowledge at (10%), enabling an a priori determination of the neutrino energy without relying on the reconstruction of final-state particles. This is the fundamental principle of the Narrow-Band Off-Axis (NBOA) technique, which can be applied to fluxes from and decays.
@@ 7 Neutrino cross-section measurements
A monitored and tagged neutrino beam offers a wide range of unique and powerful cross-section measurements that can be tailored to characterise the physics responsible for the dominant systematic uncertainties in neutrino oscillation analyses. Past and present experiments such as MiniBooNE, T2K, MicroBooNE, NOvA and MINERvA have provided a wide range of measurements on different targets. This large body of measurements, primarily focused on quasi-elastic-like interactions and in the 1 GeV regime, has highlighted important shortcomings in the way in which we model neutrino-nucleus interactions. Furthermore, in the run-up to the DUNE experiment, there are no existing or planned high-statistics measurements using an argon target in the energy range relevant for DUNE beam oscillations (i.e., 2-3 GeV and above). In this section, we give a non-exhaustive overview of some of the measurements that could be made with SBN@CERN and how they can reduce uncertainties in neutrino oscillation experiments, with a focus on their application to supporting DUNE’s physics program since we assume the use of the argon-based reference detector described in Sec. 4. Whilst future work will report tailored studies assuming a water-based detector and will project cross-section measurements focused on supporting the physics program of Hyper-Kamiokande, in this section, we give qualitative expectations of what could be achieved. The physics goals listed in this section are part of a global strategy of cross-section measurements aimed at addressing the main sources of systematic uncertainties for future experiments. They complement past measurements made by the MiniBooNE, T2K, NOvA, MicroBooNE, and MINERvA experiments, and will build on the lessons we will learn from upcoming measurements at the Fermilab SBN program with ICARUS and SBND. They are also complementary to the measurements that will be performed in-situ with the DUNE and Hyper-Kamiokande near detectors. Note that this section focuses on measurements of neutrino cross-sections, but in principle, similar measurements could be made with an antineutrino beam. Exploration of this opportunity is also left for future work.
@@
We begin in Sec. 7.1 with a description of the neutrino fluxes, and then in Sec. 7.2 we provide an overview of the simulation to estimate the event rates that can be used for cross-section measurements. We proceed to focus first on measurements that exploit the NBOA technique and the monitored neutrino beam to first make a inclusive energy-dependent cross-section measurement in Sec. 7.3 and then and differential cross-section measurements in Sec. 7.4. In Sec. 7.5, we show how the PRISM technique can be applied using the NBOA fluxes to directly measure the / cross-section ratio. We further show how the NBOA technique can be used to constrain common backgrounds for neutrino oscillation analyses, focusing on the NC contribution in Sec. 7.6. We then proceed to consider measurements benefiting from neutrino tagging in Sec. 7.7, including direct measurements of neutrino energy reconstruction bias that can be used to calibrate DUNE alongside measurements of powerful kinematic quantities to characterise neutrino interactions that are usually only accessible to electron-scattering experiments. Finally, the cross-section measurement reach of SBN@CERN is summarised in Sec. 7.8.
@@ 7.4 Flux-averaged and CC double-differential cross-section measurements
where is the square of the four momentum transfer and is the energy transfer. and are the outgoing lepton energy, momentum vector, direction with respect to the neutrino beam, and lepton mass, respectively. For this illustration, we assume that the reconstructed from particle kinematics has been unfolded to the true value of [90]. We note that this is a model-dependent procedure when using a wide neutrino beam, but the model dependence is mitigated at SBN@CERN by using tagged neutrinos, whose energy is well-known. We will describe the prospects for such measurements in Fig. 35.
@@ 7.6 Flux-averaged NC cross-section measurement
In addition to providing measurements of the CC-inclusive cross-section, SBN@CERN is also well suited to provide a measurement of neutral current (NC) processes. NC interactions constitute a source of background for neutrino oscillation measurements. Although no charged lepton is produced at the primary vertex, hadrons or photons produced in NC interactions can be mis-reconstructed as electrons following photon conversion processes. This causes some NC events to be misattributed to CC events with a final state electron. This is notably one of the main sources of background for neutrino oscillation sources, in which the flavor of the lepton is used to measure electron (anti-)neutrino appearance. The main NC channel that contributes to this background is the production of neutral pions in NC interactions, which we will refer to as the topology. The neutral pions can be produced via a resonance, coherent scattering off the nucleus, or via neutral pion production during reinteractions inside the nucleus. This process has been measured by several neutrino scattering experiments (see e.g. a measurement performed by the MicroBooNE collaboration in [96], which we use as a reference for this study). Still, the measurements are limited to energies below 1 GeV. In this section, we present the prospects of measuring this process using a monitored neutrino beam and using the NBOA technique.
@@ 7.7 Cross-section measurements with the tagged neutrino sample
In addition to the measurements using a monitored neutrino beam and the NBOA technique, SBN@CERN provides the unique opportunity to perform measurements using tagged neutrinos. This provides prospects for performing measurements where the neutrino energy is known on an event-by-event basis. In this section, we present several measurements that can be made by exploiting this new capability and discuss their impact for future generation experiments.
@@ 7.7.4 Other measurements with a tagged neutrino beam
7.8 Summary of the SBN@CERN physics potential
SBN@CERN offers the possibility to make a wide range of previously unattainable measurements that may be the key to allowing the next generation of long-baseline neutrino oscillation experiments to reach their ultimate precision. It directly confronts aspects of neutrino interaction physics expected to drive the leading sources of systematic uncertainty for neutrino oscillation experiments:
@@ 7.8 Summary of the SBN@CERN physics potential
Future work will focus on expanding these studies to include simulations using a water-based target tailored to supporting Hyper-Kamiokande’s physics program, anti-neutrino beam running, lower energy beam configurations, sensitivity to BSM physics, and a detailed assessment of the detector performance required to realize the full SBN@CERN physics program.
@@ 8 Conclusions
SBN@CERN addresses a fundamental limitation in neutrino physics: the lack of precise knowledge of neutrino cross-sections at the GeV scale. This uncertainty currently hinders the physics potential of long-baseline experiments like DUNE and Hyper-Kamiokande. It is based on two innovative techniques, whose proof-of-principle has been demonstrated in recent years by the NP06/ENUBET and NuTag Collaborations. SBN@CERN introduces a novel beam that monitors leptons at the single-particle level and can tag both the parent meson and the accompanying muon of the neutrinos interacting in the detector.
@@
While the current focus has been put on neutrino cross-sections, additional studies are being performed to address the potential of SBN@CERN in terms of BSM physics topics that could benefit from the presence of a high-granularity massive detector close to a proton and hadron beam dump associated with an unprecedented control over neutrino-originating decays. After the successful R&D phase and beam optimization, a new Collaboration is being formed to present the experiment proposal within a two-year timescale. The SBN@CERN project has the potential to revolutionize neutrino cross-section measurements, supporting the next generation of neutrino physics experiments. Leveraging CERN’s infrastructure, advanced detection techniques, and precision beam monitoring, this proposal marks a crucial step toward reducing systematic uncertainties in neutrino oscillation physics and advancing our understanding of neutrino interactions in nuclear media.

### v2

@@ 4.2 Water-based detector
Water is the target material for Hyper-Kamiokande, making a detector based on this target an invaluable tool for achieving the nuSCOPE physics goals through precise neutrino interaction measurements in water. A widely used option, water Cherenkov detectors, offer less precise reconstruction of neutrino interactions at the GeV scale compared to liquid argon detectors, but this limitation is compensated by reduced complexity and cost.
Tests of Water Cherenkov detector performance using charged beams—mirroring the ProtoDUNE physics program—are currently ongoing at CERN with the WCTE detector [59]. WCTE is a 40-ton demonstrator of the Intermediate Water Cherenkov Detector (IWCD) [60], which is currently under construction in Japan. Water Cherenkov technology offers an ideal setting for nuSCOPE operations, even in tagged mode, as it can achieve sub-nanosecond time resolution at large scales with limited light coverage, due to the inherently faster propagation of Cherenkov light compared to liquid argon scintillation light. The WCTE/IWCD technology employs multi-PMT modules instead of the classical 20-inch PMTs used in Hyper-Kamiokande, enhancing energy linearity over a broader dynamic range and thereby improving the precision of cross-section measurements. However, the 40-ton size of WCTE is insufficient to complete the full nuSCOPE physics program in water; a dedicated detector with a fiducial volume of a few hundred tons would be required. Further enhancement of WCTE’s particle identification capability could be achieved by replacing the water target with modern water-based scintillators. This novel technology, pursued by several experimental groups, aims to overcome the limitations of Water Cherenkov detectors in medium- and large-scale experiments. It is also under consideration for the aforementioned DUNE Module of Opportunity, an initiative led by the THEIA collaborations in the US and Europe.
@@ 4.3 Reference detector parameters
The studies presented here are based on a reference liquid argon (LAr) detector with a 500-ton fiducial mass, positioned 25 m from the end of the decay tunnel. The front face measures m2, with a total length of 22.3 m. The LAr TPC is assumed to have a time resolution improved to 300 ps compared to the current ProtoDUNE, thanks to an enhanced photon detection system as planned for DUNE Phase II upgrades. The final detector optimization is still a work in progress and will depend on the chosen installation site for nuSCOPE.
@@ 6.1 The Narrow-Band Off-Axis technique
Due to the proximity of the detector to the neutrino beam, and to the large size of the detector itself, the detector surface () is exposed to a range of incoming neutrino directions. The energy of produced in and decays is correlated to the emission angle, due to the two-body decay kinematics. As a result, neutrinos interacting at different off-axis angles span different energy ranges. The off-axis technique has already been exploited by a cross-section measurement from the T2K experiment [66], and is central for the PRISM technique [67], which is planned for use in the DUNE [68] and Hyper-Kamiokande near detectors [4] as well as in the SBND experiment [20]. A key advantage of a narrow-band beam like nuSCOPE is that the off-axis neutrino energy estimation can leverage the parent meson momentum knowledge at (10%), enabling an a priori determination of the neutrino energy without relying on the reconstruction of final-state particles. This is the fundamental principle of the Narrow-Band Off-Axis (NBOA) technique, which can be applied to fluxes from and decays.
@@ 7 Neutrino cross-section measurements
A monitored and tagged neutrino beam offers a wide range of unique and powerful cross-section measurements that can be tailored to characterise the physics responsible for the dominant systematic uncertainties in neutrino oscillation analyses. Past and present experiments such as MiniBooNE, T2K, MicroBooNE, NOvA and MINERvA have provided a wide range of measurements on different targets. This large body of measurements, primarily focused on quasi-elastic-like interactions and in the 1 GeV regime, has highlighted important shortcomings in the way in which we model neutrino-nucleus interactions. Furthermore, in the run-up to the DUNE experiment, there are no existing or planned high-statistics measurements using an argon target in the energy range relevant for DUNE beam oscillations (i.e., 2-3 GeV and above). In this section, we give a non-exhaustive overview of some of the measurements that could be made with nuSCOPE and how they can reduce uncertainties in neutrino oscillation experiments, with a focus on their application to supporting DUNE’s physics program since we assume the use of the argon-based reference detector described in Sec. 4. Whilst future work will report tailored studies assuming a water-based detector and will project cross-section measurements focused on supporting the physics program of Hyper-Kamiokande, in this section, we give qualitative expectations of what could be achieved. The physics goals listed in this section are part of a global strategy of cross-section measurements aimed at addressing the main sources of systematic uncertainties for future experiments. They complement past measurements made by the MiniBooNE, T2K, NOvA, MicroBooNE, and MINERvA experiments, and will build on the lessons we will learn from upcoming measurements at the Fermilab SBN program with ICARUS and SBND. They are also complementary to the measurements that will be performed in-situ with the DUNE and Hyper-Kamiokande near detectors. Note that this section focuses on measurements of neutrino cross-sections, but in principle, similar measurements could be made with an antineutrino beam. Exploration of this opportunity is also left for future work.
@@
We begin in Sec. 7.1 with a description of the neutrino fluxes, and then in Sec. 7.2 we provide an overview of the simulation to estimate the event rates that can be used for cross-section measurements. We proceed to focus first on measurements that exploit the NBOA technique and the monitored neutrino beam to first make a inclusive energy-dependent cross-section measurement in Sec. 7.3 and then and differential cross-section measurements in Sec. 7.4. In Sec. 7.5, we show how the PRISM technique can be applied using the NBOA fluxes to directly measure the / cross-section ratio. We further show how the NBOA technique can be used to constrain common backgrounds for neutrino oscillation analyses, focusing on the NC contribution in Sec. 7.6. We then proceed to consider measurements benefiting from neutrino tagging in Sec. 7.7, including direct measurements of neutrino energy reconstruction bias that can be used to calibrate DUNE alongside measurements of powerful kinematic quantities to characterise neutrino interactions that are usually only accessible to electron-scattering experiments. Finally, the cross-section measurement reach of nuSCOPE is summarised in Sec. 7.8.
@@ 7.4 Flux-averaged and CC double-differential cross-section measurements
where is the square of the four momentum transfer and is the energy transfer. and are the outgoing lepton energy, momentum vector, direction with respect to the neutrino beam, and lepton mass, respectively. For this illustration, we assume that the reconstructed from particle kinematics has been unfolded to the true value of [90]. We note that this is a model-dependent procedure when using a wide neutrino beam, but the model dependence is mitigated at nuSCOPE by using tagged neutrinos, whose energy is well-known. We will describe the prospects for such measurements in Fig. 35.
@@ 7.6 Flux-averaged NC cross-section measurement
In addition to providing measurements of the CC-inclusive cross-section, nuSCOPE is also well suited to provide a measurement of neutral current (NC) processes. NC interactions constitute a source of background for neutrino oscillation measurements. Although no charged lepton is produced at the primary vertex, hadrons or photons produced in NC interactions can be mis-reconstructed as electrons following photon conversion processes. This causes some NC events to be misattributed to CC events with a final state electron. This is notably one of the main sources of background for neutrino oscillation sources, in which the flavor of the lepton is used to measure electron (anti-)neutrino appearance. The main NC channel that contributes to this background is the production of neutral pions in NC interactions, which we will refer to as the topology. The neutral pions can be produced via a resonance, coherent scattering off the nucleus, or via neutral pion production during reinteractions inside the nucleus. This process has been measured by several neutrino scattering experiments (see e.g. a measurement performed by the MicroBooNE collaboration in [96], which we use as a reference for this study). Still, the measurements are limited to energies below 1 GeV. In this section, we present the prospects of measuring this process using a monitored neutrino beam and using the NBOA technique.
@@ 7.7 Cross-section measurements with the tagged neutrino sample
In addition to the measurements using a monitored neutrino beam and the NBOA technique, nuSCOPE provides the unique opportunity to perform measurements using tagged neutrinos. This provides prospects for performing measurements where the neutrino energy is known on an event-by-event basis. In this section, we present several measurements that can be made by exploiting this new capability and discuss their impact for future generation experiments.
@@ 7.7.4 Other measurements with a tagged neutrino beam
7.8 Summary of the nuSCOPE physics potential
nuSCOPE offers the possibility to make a wide range of previously unattainable measurements that may be the key to allowing the next generation of long-baseline neutrino oscillation experiments to reach their ultimate precision. It directly confronts aspects of neutrino interaction physics expected to drive the leading sources of systematic uncertainty for neutrino oscillation experiments:
@@ 7.8 Summary of the nuSCOPE physics potential
Future work will focus on expanding these studies to include simulations using a water-based target tailored to supporting Hyper-Kamiokande’s physics program, anti-neutrino beam running, lower energy beam configurations, sensitivity to BSM physics, and a detailed assessment of the detector performance required to realize the full nuSCOPE physics program.
@@ 8 Conclusions
nuSCOPE addresses a fundamental limitation in neutrino physics: the lack of precise knowledge of neutrino cross-sections at the GeV scale. This uncertainty currently hinders the physics potential of long-baseline experiments like DUNE and Hyper-Kamiokande. It is based on two innovative techniques, whose proof-of-principle has been demonstrated in recent years by the NP06/ENUBET and NuTag Collaborations. nuSCOPE introduces a novel beam that monitors leptons at the single-particle level and can tag both the parent meson and the accompanying muon of the neutrinos interacting in the detector.
@@
While the current focus has been put on neutrino cross-sections, additional studies are being performed to address the potential of nuSCOPE in terms of BSM physics topics that could benefit from the presence of a high-granularity massive detector close to a proton and hadron beam dump associated with an unprecedented control over neutrino-originating decays. After the successful R&D phase and beam optimization, a new Collaboration is being formed to present the experiment proposal within a two-year timescale. The nuSCOPE project has the potential to revolutionize neutrino cross-section measurements, supporting the next generation of neutrino physics experiments. Leveraging CERN’s infrastructure, advanced detection techniques, and precision beam monitoring, this proposal marks a crucial step toward reducing systematic uncertainties in neutrino oscillation physics and advancing our understanding of neutrino interactions in nuclear media.